OpenStudy (anonymous):

Trig question on double angle formulas: Sin theta=3/5, 0

OpenStudy (fibonaccichick666):

well, actually 3/5 is a special angle isn't it?

OpenStudy (anonymous):

is it?

OpenStudy (aum):



OpenStudy (fibonaccichick666):

I'm pretty sure, maybe you could draw a picture to think about it

OpenStudy (fibonaccichick666):

or that^

OpenStudy (aum):

a) sin(2t) = 2sin(t)cos(t) sin(t) = 3/5; cos(t) = 4/5 sin(2t) = 2 * 3/5 * 4/5 = 24/25

OpenStudy (anonymous):

ok I have that part down, and the answers for a and b. I have sin (2 theta)=-24/25 and cos (2 theta)=7/25. I'm not really sure how the formulas work on c and d though the whole thing has me a bit confused

OpenStudy (fibonaccichick666):

How did you get negative?

OpenStudy (fibonaccichick666):

let's look at the first two before we get to C and D

OpenStudy (anonymous):

I got a positive but the book told me it was a negative and that x was less than zero but I don't know were that came from, it wasn't a part of the original problem